# TempoControl: Temporal Attention Guidance for Text-to-Video Models

### v1

@@ 5 Experiments
Below, we describe the applications and evaluation protocols used in our experiments.
Setup.
We use Wan 2.1 1.3B as our text-to-video backbone. All experiments were run on a single NVIDIA H200 GPU. TempoControl takes approximately $\sim$460s per video, compared to $\sim$170s for vanilla generation. Hyperparameter details and ablations are available in the appendix.
@@
We employ Wan 2.1 1.3B with explicit temporal cues in the prompt (e.g., “…at the third second…”). We describe the exact phrasing details in the appendix.
@@
We evaluate on VBench (Huang et al. 2024), reporting temporal-related metrics: Subject Consistency (main subject coherence across frames), Background Consistency (background stability), Motion Smoothness (absence of temporal flicker), and Dynamic Degree (overall motion magnitude), as well as quality metrics: Aesthetic Quality (visual appeal) and Imaging Quality (clarity of frames). Additional details about these metrics are provided in the appendix.
We also introduce a new metric, Temporal Accuracy, which measures whether objects appear in the correct frames as specified by the prompt, and provides breakdowns for both object absence and presence. By default, the official VBench metric uses the GRiT detector (Wu et al. 2024). To provide a more up-to-date assessment, we report temporal accuracy results using YOLOv10 (Wang et al. 2024a).
For movement, Temporal Accuracy is assessed by matching the movement time specified in the prompt with the second exhibiting the most movement, as determined by optical flow. Each video is sampled at 4 frames per second; we compute the average flow magnitude per second, and the second with the highest average is considered the most active.
@@ 5.1 Quantitative Results
In Table 1 (One Object), our method improves temporal accuracy by 17.06% to 81.00%, with similar gains in temporal absence (+21.37) and presence (+12.41) accuracy compared to text prompts with explicit time cues. Image Quality and Aesthetic Quality remain comparable, with a difference of 1%.
We observe decreases in Dynamic Degree, dropping by 8.75% to 38.75%, reflecting reduced dynamics due to temporal control, which requires object presence over extended periods of the video.
@@
In Table 1 (Two Objects), we show that the two-object setting presents a significant challenge, with the baseline achieving only 37.5% temporal accuracy. Our method improves to 55.0%. Temporal absence and presence accuracies also show gains of 20.6% and 14.4%, respectively. Imaging and Aesthetic Quality increase by about 2% each. Subject Consistency, Background Consistency, and Motion Smoothness remain similar between methods. At the same time, we find that Dynamic Degree is reduced here, as in the one-object case, by roughly 21%. This reflects the trade-off imposed by temporal control, which requires static background objects and reduces dynamic content.
Figure 4: Examples of videos generated by Wan 2.1 (left) and our method with temporal conditioning (right) for prompts that specify temporal conditions. The prompts are simplified for brevity. We show frames corresponding to the first, middle, and last third of each video.
@@
Movement temporal control.
Table 1 (Movement) shows that TempoControl significantly improves temporal accuracy (53% vs. 19%) compared to the text-prompt baseline. We note that the original temporal accuracy of the model is similar to random, indicating that Wan 2.1 does not capture movement cues at all. The Dynamic Degree here increases, as the goal is to introduce more movement, demonstrating that our approach successfully creates more pronounced motion. Imaging and Aesthetic Quality remain on par with the baseline. Overall, our explicit temporal guidance sharpens movement timing without compromising visual quality and consistency.
Multiple objects benchmark.
We also report results on the VBench Multiple Objects benchmark in Table 2, which tests the ability to generate two objects described in the input text throughout the entire video. We observe improved performance on the Multiple Object metric (76.34% vs. 74.13%), although this comes at the cost of reduced video dynamics. Detailed results are provided in the appendix.
Human evaluation.
We conducted a human evaluation with 50 annotators, each comparing 16 videos (8 one-object, 8 two-object) generated by TempoControl and Wan 2.1. Annotators selected which video was better in terms of temporal accuracy and visual quality, or if they were the same. As shown in Table 3, our method outperformed the baseline in temporal accuracy (61.51% vs. 16.94%) and also improved visual quality. For temporal accuracy, 21.55% of responses indicated no preference, likely due to subtle timing differences. Full details of the evaluation are provided in the appendix.
@@ 5.2 Qualitative Results
In Figure 4, we show four videos generated to depict objects or actions at specific times. In the first row, where a single apple is expected to appear at the fourth second, Wan 2.1 generates the apple near the middle of the five-second video, while our method correctly places it near the end of the video. For the “cat and dog, dog at first half” prompt, Wan 2.1 fails to show the dog at the beginning of the video, whereas our approach accurately places the dog in the first half. In the “magician pulling a rabbit last second” scenario, Wan 2.1 displays the rabbit from the start, instead of only in the final second as required. Similarly, for “tiger leaping first second,” Wan 2.1 does not capture the leaping motion at the correct time, while our method, with temporal conditioning, accurately generates the requested action. All videos are included in the supplementary material.
@@ 5.3 Audio-Visual Alignment
We also experimented with zero-shot alignment of video to external audio by using a preprocessed audio envelope as the temporal condition. Given a waveform $y$, we extract its onset strength envelope $s_{t}$, adaptively pool it to match the number of frames, and normalize to $[0,1]$ via min-max normalization. Prominent peaks are preserved via thresholding and Gaussian smoothing:
$\tilde{s}_{t}=\begin{cases}\hat{s}_{t},&\text{if }\hat{s}_{t}\geq\tau_{audio},\\ \text{GaussianSmooth}(\hat{s}_{t};\sigma),&\text{otherwise}.\end{cases}$ (11)
where $\tau_{audio}$ is a threshold for peak preservation, and $\sigma$ is the standard deviation of the Gaussian kernel used for smoothing.
In Figure 5, we present examples of accurate video adaptation to audio peaks, with additional results in the appendix. Although experiments are limited to simple cases with prominent peaks, these findings suggest a promising direction for aligning generated video with audio without requiring costly training on curated datasets.
5.4 Discussions
Text temporal control limitations.
In our experiments, adding explicit timing to Wan 2.1 prompts did not improve temporal accuracy and often reduced video and image quality. For example, temporal accuracy fell from 65% to 64%, and video quality from 56% to 53%. For fairness, all reported results use the same prompt setup as the text baseline, despite omitting explicit timing, which yields higher performance for our approach. For instance, for a single object, we observe a video quality of 56% compared to 53%, with the same temporal accuracy of 81%. See the appendix for details.
Consistency limitations.
Temporal conditioning may lead to perceptual changes in background or unspecified features. For example, it can alter the color of a cat (see Figure 1). Although perceptual consistency is often preserved, our current objectives do not fully guarantee it, which remains a promising area for future work.
@@ 5.5 Conclusions
Supplementary Material: TempoControl – Temporal Attention Guidance for Text-to-Video Models
This supplementary material provides detailed information supporting our main paper. It includes algorithmic implementation, experimental design, application-specific configurations, extended quantitative and qualitative results, and a human evaluation setup. Videos and code are provided as part of the supplementary materials package.
Specifically, we include:
@@
    Full algorithm and pseudocode (Section .6).
@@
.6 Algorithm Pseudo-code
@@
The optimization loss $\mathcal{L}$ combines four components: correlation alignment $\mathcal{L}_{\text{c}}^{t}$, energy modulation $\mathcal{L}_{\text{enr}}^{t}$, entropy regularization $\mathcal{L}_{\text{ent}}^{t}$, and spatial consistency $\mathcal{L}_{\text{s}}^{t}$. The latent code is updated via gradient descent for up to $l$ steps per denoising timestep, or until the Pearson correlation exceeds a threshold $\tau_{\text{corr}}$. After optimization, the standard denoising process resumes to produce the final latent $z_{0}$.
@@
12:     $\mathcal{L}_{\text{enr}}^{t}\leftarrow\mathcal{L}_{\ominus}^{t}-\mathcal{L}_{\oplus}^{t}$
@@
14:     $\mathcal{L}_{\text{s}}^{t}\leftarrow\sum_{j=1}^{T^{\prime}}\|\bar{A}^{T}_{j,i}-\bar{A}^{t}_{j,i}\|_{2}^{2}$
15:     $\mathcal{L}\leftarrow\mathcal{L}+\mathcal{L}_{\text{c}}^{t}+\lambda_{1}\mathcal{L}_{\text{enr}}^{t}+\lambda_{2}\mathcal{L}_{\text{ent}}^{t}+\lambda_{3}\mathcal{L}_{\text{s}}^{t}$
16:   end for
17:   $z_{t}^{\prime}\leftarrow z_{t}-\alpha_{t}\nabla_{z_{t}}(\mathcal{L}/N)$
18:   if All $\mathcal{L}_{\text{corr}}^{t}\leq\tau_{\text{corr}}$ then
19:     break
20:   end if
21:   $z_{t}\leftarrow z_{t}^{\prime}$
22:  end for
23:  $z_{t-1}\leftarrow\operatorname{DiT}(z_{t}^{\prime},\mathcal{P},t)$
24:end for
25:Continue denoising using the standard denoising diffusion process for $t=T{-}k,\dots,1$
26:return $z_{0}$
.7 Experimental Setup
Hyperparameters were tuned early using a small validation set specific to each task. While large-scale ablations were limited by computational constraints, we found the method to be robust across scenarios with only minor adjustments.
All experiments are conducted using Wan 2.1 as the backbone. We apply a classifier-free guidance scale of 6 and a sample shift of 3 during generation.
For single-object scenes, we use $\lambda_{1}{=}1$, $\lambda_{2}{=}10$, $\lambda_{3}{=}1$, a learning rate of $0.0005$, a Pearson stopping threshold of $\tau_{\text{corr}}{=}0.9$, and apply optimization for the first l=10l{=}10 denoising steps with up to k=10k{=}10 gradient updates per step.
In two-object scenes, which require more aggressive guidance, we raise the learning rate to $0.001$ for improved convergence. For motion-centric prompts, we reduce the temporal alignment weight to $\lambda_{1}{=}0.3$ and relax the Pearson threshold to $\tau_{\text{corr}}{=}0.85$ to allow more flexible optimization.
.8 Applications Setup
@@
To assess temporal grounding for individual objects, we construct a dataset using the 80 object categories detectable by the YOLO detector. These span common classes, such as animals and vehicles, as well as specific items, including handbags and sports equipment. Videos are generated from prompts of the form: “An empty scene. Suddenly, $\langle o_{1}\rangle$ appears out of nowhere, drawing all attention.” Object classes are uniformly distributed across different time intervals.
@@
We evaluate TempoControl on 82 object pairs from the VBench protocol. Prompts take the form: “The video begins with a serene view centered on $\langle o_{1}\rangle$, with no sign of $\langle o_{2}\rangle$. In the second half, $\langle o_{2}\rangle$ unexpectedly appears, altering the dynamic of the scene.” A static control signal (all ones) is used for $\langle o_{1}\rangle$. We sample 20 evenly spaced frames and require $\langle o_{1}\rangle$ to always be present. A frame is counted as successful if both objects are detected when $\langle o_{2}\rangle$ is active, and only $\langle o_{1}\rangle$ is detected when $\langle o_{2}\rangle$ is inactive.

### v2

@@ 5 Experiments
Experimental setup.
All experiments are conducted on a single NVIDIA H200 GPU using Wan 2.1-S as the backbone. Due to limited resources, we focus on the smaller model, which already outperforms larger baselines. We optimize the latent representation during the first 10 denoising steps, applying up to 10 gradient updates per step with a learning rate of $5\times 10^{-4}$, increased to $1\times 10^{-3}$ for the two-object setting. We use $\lambda_{1}{=}0.3$ and $\lambda_{2}{=}10$. See the appendix for ablation details.
@@
We evaluate our method on several text-to-video backbones, including Wan 2.1 (1.3B and 14B), Wan 2.2-TI2V (5B), and CogX1.5 (5B). As a baseline, we use each model with explicit temporal cues in the prompt (e.g., “…at the third second …”). The exact phrasing templates are provided in the appendix. Note that we also use the same timing templates in our approach for fair comparison. The lack of dedicated baselines stems from the novelty of our objective, which involves fine-grained temporal control and necessitates the definition of new evaluation metrics, discussed next.
@@
Current benchmarks such as VBench [14] do not measure temporal control. To address this, we introduce Temporal Accuracy, which checks whether each token appears in the correct frames specified by the prompt, with separate scores for presence and absence. VBench relies on the GRiT detector [29], which we find unreliable for small or partially visible objects. Instead, we detect objects using the more recent YOLOv10 [24].
For motion-centric prompts, Temporal Accuracy aligns the prompt-specified action time with the second exhibiting the strongest motion, as determined by per-second optical flow magnitude computed from videos sampled at 4 fps. Full details are provided in the appendix.
We also report results on the VBench Multiple Object benchmark (detecting object appearance throughout the video), and the Imaging Quality metric that uses a perceptual image-quality model.
Benchmark.
In the single-object setup, we utilize 80 YOLOv10 classes, which span animals, everyday objects, vehicles, and more. For the two-object setting, we use 82 object pairs based on the VBench multiple-object benchmark. For movement, we curate 100 types of actions. Temporal masks are generated together with the prompt templates to ensure alignment. Full prompt patterns and categories are provided in the appendix.
Table 1: Comparison across one-object, two-object, and movement temporal-control tasks. Baselines use text-only temporal cues, while Ours applies TempoControl on Wan2.1 (1.3B). Bold values indicate the best result in each block.
Method # Temporal Accuracy T. Absence Accuracy T. Presence Accuracy Imaging Quality
One Object
CogX1.5 5B 61.43% 48.64% 75.50% 65.36%
Wan2.1-S 1.3B 63.94% 67.38% 60.50% 53.76%
Wan2.1-L 14B 64.23% 55.34% 74.00% 66.89%
Wan2.2 5B 78.02% 87.42% 68.00% 48.13%
Ours 1.3B 82.50% 83.25% 81.75% 56.51%
Two Objects
CogX1.5 5B 24.68% 46.46% 4.88% 63.78%
Wan2.1-S 1.3B 37.50% 45.85% 29.15% 68.56%
Wan2.1-L 14B 40.07% 54.15% 27.27% 69.01%
Wan2.2 5B 47.05% 56.86% 36.59% 67.83%
Ours 1.3B 52.93% 58.05% 47.80% 71.09%
Movement
CogX1.5 5B 21% 61.39%
Wan2.1-S 1.3B 19% 60.46%
Wan2.1-L 14B 18% 61.31%
Wan2.2 5B 23% 60.61%
Ours 1.3B 57% 63.38%
@@ 5.1 Quantitative Results
In Table 1 (One Object), our method improves temporal accuracy by 18.56%, with gains of 15.87% in absence accuracy and 21.25% in presence accuracy over the Wan2.1-S baseline. Imaging Quality also increases by about 3%.
Larger baseline models achieve higher imaging quality, as expected, yet all fail to provide reliable temporal control from text alone. Wan2.2’s strong absence accuracy largely reflects degraded image quality, which causes detectors to miss objects. In a separate experiment, we tested removing the timing cues, which increased its Imaging Quality to 57.78%, indicating that it is not robust to temporal instructions; details are in the appendix.
@@
In Table 1 (Two Objects), the two-object setting is notably challenging. Wan2.1-S baseline achieves only 37.50% temporal accuracy, while our method improves this to 52.93% (+15.43). Absence and presence accuracies also increase by 12.20% and 18.65%, respectively, and Imaging Quality improves by 2.53%.
Larger models also show similar degradation in temporal control in the two-object setting. Conversely, Imaging Quality is consistently higher across all methods, suggesting a possible bias in the perceptual metric toward multi-object scenes. Our method achieves the highest Imaging Quality.
Movement temporal control.
Table 1 (Movement) evaluates temporal accuracy when conditioning is indicative of a peak of action. The temporal accuracy measures the timing of when the action peaks (see Metrics). Wan2.1-S baseline achieves only 19% temporal accuracy, indicating that the model does not follow action timing from text alone. TempoControl improves this to 57%, showing a substantial gain in aligning motion onset with the prompt.
Low performance in movement conditioning is consistent across all models. Our Imaging Quality remains the best in this setup as well.
Multiple objects benchmark.
We also report results on the VBench Multiple Objects benchmark in Table 2, which tests the ability to generate two objects described in the input text throughout the entire video. We observe improved performance on the Multiple Object metric (76.34% vs. 74.13%). Detailed results are provided in the appendix.
Ablation.
Table 3 shows that each component of our objective contributes differently to temporal control and visual quality. Pearson correlation (C) provides the strongest temporal alignment, driven largely by improvements in T. Absence Accuracy, but it also reduces Imaging Quality. Closer inspection suggests that degraded visuals can cause detectors to miss objects, artificially inflating absence scores. Entropy regularization (E) improves presence accuracy and stabilizes spatial attention. Combining both (C+E) yields balanced gains across metrics. Notably, using entropy alone produces the highest Imaging Quality (59.52%), outperforming even the text-only baseline, suggesting that entropy-based attention smoothing is broadly beneficial for video generation. Additional ablations are provided in the appendix.
Human evaluation.
We conducted a user study with 50 computer science graduate students, each comparing 16 video pairs (8 one-object and 8 two-object) generated by TempoControl and Wan 2.1 with text-based temporal cues. Annotators judged temporal accuracy and visual quality, or indicated no difference, which we asked them to use sparingly (full protocol details are provided in the appendix). As shown in Table 4, our method is preferred for temporal accuracy (61.51% vs. 16.94%) and also improves visual quality. Notably, 21.55% of temporal comparisons reported no preference, reflecting that humans struggle to perceive minor timing differences.
Table 2: Full results for the multiple-object benchmark. Best results per column are in bold.
Method Multiple Object (GRiT) Multiple Object (YOLO) Imaging Quality
Text 74.13% 61.54% 70.25%
Ours 76.37% 65.73% 70.21%
Table 3: Ablation results for the one-object setting. C = Pearson correlation, E = Entropy.
Method Temporal Accuracy T. Absence Accuracy T. Presence Accuracy Imaging Quality
Text 63.94% 67.38% 60.50% 53.76%
Only C 81.19% 91.50% 70.88% 50.96%
Only E 72.94% 66.88% 79.00% 59.52%
C + E 78.38% 77.38% 79.38% 57.60%
Ours 82.50% 83.25% 81.75% 56.51%
Table 4: User study comparing our method with Wan2.1 in terms of temporal accuracy and visual quality.
Question Type Ours Baseline Same
Temporal Accuracy 61.51% 16.94% 21.55%
Visual Quality 62.66% 25.99% 11.35%
Figure 5: Examples generated by Wan 2.1 (left) and our temporally conditioned method (right) for prompts with timing constraints. Prompts are simplified for brevity. Shown are frames from the first, middle, and final thirds of each video.
@@ 5.2 Qualitative Results
In Figure 5, each row illustrates a different temporal-control task. The first row shows a one-object example: for the prompt “an apple appears at the fourth second,” Wan 2.1 places the apple near the middle of the five-second video, whereas our method correctly places it near the end. The second row shows a two-object example: for “a car and a bicycle, car in the second half,” Wan 2.1 displays the car at the beginning instead of the specified time. The third and fourth rows depict movement-based prompts. For a magician pulling a rabbit at the last second, Wan 2.1 shows the rabbit throughout the video, whereas our method reveals it only at the end. For a tiger leaping in the first second, Wan 2.1 fails to time the leap, whereas our method accurately aligns the motion with the prompt. Additional examples are included in the supplementary material.
@@ 5.3 Audio-Visual Alignment
We also explore zero-shot alignment of video to external audio by using a preprocessed audio envelope as the temporal condition. Given a waveform $y$, we compute its onset-strength envelope $s_{t}$, pool it to the number of frames, and normalize it to $[0,1]$. Peaks are preserved through thresholding and otherwise smoothed with a Gaussian filter:
$\tilde{s}_{t}=\begin{cases}\hat{s}_{t},&\hat{s}_{t}\geq\tau_{audio},\\ \text{GaussianSmooth}(\hat{s}_{t};\sigma),&\text{otherwise}.\end{cases}$ (10)
Figure 6 shows examples where video motion aligns with peaks in the audio envelope, with additional results in the appendix. While tested on simple cases with clear peaks, these results suggest that audio-driven temporal conditioning is a promising approach for achieving audio-video alignment without requiring training on paired datasets.
Limitations.
TempoControl has two main limitations. First, it increases inference time to approximately 460 seconds per video, compared to 170 seconds for vanilla generation. Nonetheless, it remains an inference-only approach that requires no additional supervision, thereby avoiding the degradation in robustness and diversity often seen in supervised training. Second, our conditioning can introduce mild perceptual shifts in target attributes (for example, altering the color of a cat; see Figure 1). While consistency is often maintained, our current objectives do not fully guarantee it, making this an interesting direction for future work.
Temporal control in text-to-video models.
We observe that adding explicit timing cues to current text-to-video models does not improve temporal accuracy and often degrades visual quality. For example, in Wan 2.1 (one-object setting), temporal accuracy remains at 64% while Imaging Quality drops from 56% to 53%. This highlights a broader limitation: existing models lack reliable temporal grounding from text alone. For fairness, we keep the same prompt format for all baselines, though omitting explicit timing slightly improves our Imaging Quality (56% vs. 53%) without affecting Temporal Accuracy (details in the appendix).
@@ 5.4 Conclusions
Supplementary Material for TempoControl: Temporal Attention Guidance for Text-to-Video Models
This supplementary material includes algorithmic pseudocode, ablation studies, experimental design details, application-specific configurations, extended quantitative and qualitative results, and the human-evaluation setup. Videos and code are also provided in the supplementary package.
The sections of the supplementary material are as follows:
@@
    Algorithm pseudocode (Section .5).
@@
.5 Algorithm Pseudo-code
@@
The optimization loss $\mathcal{L}$ combines three components: correlation alignment $\mathcal{L}_{\text{c}}^{t}$, magnitude modulation $\mathcal{L}_{\text{enr}}^{t}$, entropy regularization and $\mathcal{L}_{\text{ent}}^{t}$. The latent code is updated via gradient descent for up to $l$ steps per denoising timestep, or until the Pearson correlation exceeds a threshold $\tau_{\text{corr}}$. After optimization, the standard denoising process resumes to produce the final latent $z_{0}$.
@@
12:     $\mathcal{L}_{\text{mag}}^{t}\leftarrow\mathcal{L}_{\ominus}^{t}-\mathcal{L}_{\oplus}^{t}$
@@
14:     $\mathcal{L}\leftarrow\mathcal{L}+\mathcal{L}_{\text{c}}^{t}+\lambda_{1}\mathcal{L}_{\text{enr}}^{t}+\lambda_{2}\mathcal{L}_{\text{ent}}^{t}$
15:   end for
16:   $z_{t}^{\prime}\leftarrow z_{t}-\alpha_{t}\nabla_{z_{t}}(\mathcal{L}/N)$
17:   if All $\mathcal{L}_{\text{corr}}^{t}\leq\tau_{\text{corr}}$ then
18:     break
19:   end if
20:   $z_{t}\leftarrow z_{t}^{\prime}$
21:  end for
22:  $z_{t-1}\leftarrow\operatorname{DiT}(z_{t}^{\prime},\mathcal{P},t)$
23:end for
24:Continue denoising using the standard denoising diffusion process for $t=T{-}k,\dots,1$
25:return $z_{0}$
.6 Experimental Setup
Hyperparameters were tuned early on using a small validation split tailored to each task. Although large-scale ablations were constrained by computational cost, the method proved robust across settings, requiring only minor adjustments.
All experiments use Wan 2.1 as the backbone model. We apply a classifier-free guidance scale of 6 and a sample shift of 3 during generation.
For single-object scenes, we use $\lambda_{1}{=}0.3$, $\lambda_{2}{=}10$, a learning rate of $5{\times}10^{-4}$, a Pearson stopping threshold of $\tau_{\text{corr}}{=}0.9$, and perform optimization over the first l=10l{=}10 denoising steps with up to k=10k{=}10 gradient updates per step.
For two-object scenes, which demand stronger guidance, we increase the learning rate to $10^{-3}$ to improve convergence. For motion-centric prompts, we relax the Pearson threshold to $\tau_{\text{corr}}{=}0.85$ to enable more flexible optimization.
.7 Applications Setup
@@
To evaluate temporal grounding for individual objects, we construct a dataset based on the 80 object categories detectable by the YOLO detector. These categories cover common entities such as animals, vehicles, household items, and everyday objects. The complete class list is provided below:
person, bicycle, car, motorcycle, airplane, bus, train, truck, boat, traffic light, fire hydrant, stop sign, parking meter, bench, bird, cat, dog, horse, sheep, cow, elephant, bear, zebra, giraffe, backpack, umbrella, handbag, tie, suitcase, frisbee, skis, snowboard, sports ball, kite, baseball bat, baseball glove, skateboard, surfboard, tennis racket, bottle, wine glass, cup, fork, knife, spoon, bowl, banana, apple, sandwich, orange, broccoli, carrot, hot dog, pizza, donut, cake, chair, couch, potted plant, bed, dining table, toilet, tv, laptop, mouse, remote, keyboard, cell phone, microwave, oven, toaster, sink, refrigerator, book, clock, vase, scissors, teddy bear, hair drier, toothbrush.
Videos are generated from prompts of the form ‘An empty scene. Suddenly, $\langle o_{1}\rangle$ appears out of nowhere, drawing all attention.’ Object classes are uniformly sampled across different temporal intervals.
@@
We evaluate TempoControl on 82 prompts derived from the VBench protocol, each consisting of a pair of objects. The complete numbered list of pairs is provided below:
  1. 1.
    a bird and a cat
  2. 2.
    a cat and a dog
  3. 3.
    a dog and a horse
  4. 4.
    a horse and a sheep
  5. 5.
    a sheep and a cow
  6. 6.
    a cow and an elephant
  7. 7.
    an elephant and a bear
  8. 8.
    a bear and a zebra
  9. 9.
    a zebra and a giraffe
  10. 10.
    a giraffe and a bird
  11. 11.
    a chair and a couch
  12. 12.
    a couch and a potted plant
  13. 13.
    a potted plant and a tv
  14. 14.
    a tv and a laptop
  15. 15.
    a laptop and a remote
  16. 16.
    a remote and a keyboard
  17. 17.
    a keyboard and a cell phone
  18. 18.
    a cell phone and a book
  19. 19.
    a book and a clock
  20. 20.
    a clock and a backpack
  21. 21.
    a backpack and an umbrella
  22. 22.
    an umbrella and a handbag
  23. 23.
    a handbag and a tie
  24. 24.
    a tie and a suitcase
  25. 25.
    a suitcase and a vase
  26. 26.
    a vase and scissors
  27. 27.
    scissors and a teddy bear
  28. 28.
    a teddy bear and a frisbee
  29. 29.
    a frisbee and skis
  30. 30.
    skis and a snowboard
  31. 31.
    a snowboard and a sports ball
  32. 32.
    a sports ball and a kite
  33. 33.
    a kite and a baseball bat
  34. 34.
    a baseball bat and a baseball glove
  35. 35.
    a baseball glove and a skateboard
  36. 36.
    a skateboard and a surfboard
  37. 37.
    a surfboard and a tennis racket
  38. 38.
    a tennis racket and a bottle
  39. 39.
    a bottle and a chair
  40. 40.
    an airplane and a train
  41. 41.
    a train and a boat
  42. 42.
    a boat and an airplane
  43. 43.
    a bicycle and a car
  44. 44.
    a car and a motorcycle
  45. 45.
    a motorcycle and a bus
  46. 46.
    a bus and a traffic light
  47. 47.
    a traffic light and a fire hydrant
  48. 48.
    a fire hydrant and a stop sign
  49. 49.
    a stop sign and a parking meter
  50. 50.
    a parking meter and a truck
  51. 51.
    a truck and a bicycle
  52. 52.
    a toilet and a hair drier
  53. 53.
    a hair drier and a toothbrush
  54. 54.
    a toothbrush and a sink
  55. 55.
    a sink and a toilet
  56. 56.
    a wine glass and a chair
  57. 57.
    a cup and a couch
  58. 58.
    a fork and a potted plant
  59. 59.
    a knife and a tv
  60. 60.
    a spoon and a laptop
  61. 61.
    a bowl and a remote
  62. 62.
    a banana and a keyboard
  63. 63.
    an apple and a cell phone
  64. 64.
    a sandwich and a book
  65. 65.
    an orange and a clock
  66. 66.
    broccoli and a backpack
  67. 67.
    a carrot and an umbrella
  68. 68.
    a hot dog and a handbag
  69. 69.
    a pizza and a tie
  70. 70.
    a donut and a suitcase
  71. 71.
    a cake and a vase
  72. 72.
    an oven and scissors
  73. 73.
    a toaster and a teddy bear
  74. 74.
    a microwave and a frisbee
  75. 75.
    a refrigerator and skis
  76. 76.
    a bicycle and an airplane
  77. 77.
    a car and a train
  78. 78.
    a motorcycle and a boat
  79. 79.
    a person and a toilet
  80. 80.
    a person and a hair drier
  81. 81.
    a person and a toothbrush
  82. 82.
    a person and a sink
Prompts follow the structure: ‘The video begins with a serene view centered on $\langle o_{1}\rangle$, with no sign of $\langle o_{2}\rangle$. In the second half, $\langle o_{2}\rangle$ unexpectedly appears, altering the dynamic of the scene.’
A static control signal (all ones) is used for $\langle o_{1}\rangle$. We sample 20 evenly spaced frames and require $\langle o_{1}\rangle$ to be present throughout. A frame is counted as successful if both objects are detected when $\langle o_{2}\rangle$ is active, and only $\langle o_{1}\rangle$ is detected when $\langle o_{2}\rangle$ is inactive.